OpenStudy (anonymous):

What is the mass, in grams, of a sample of 6.98 × 1024 atoms of magnesium (Mg)?

OpenStudy (anonymous):

1mol Mg contains 6.022*10^23 atoms 6.98*10^24 atoms = (6.98*10^24)/6.022*10^23) = 11.59mol Molar mass Mg = 24.305g/mol 11.59mol = 11.59*24.305 = 281.72g Mg in sample.

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

whar kind of science classes are you in? im in chemistry

OpenStudy (anonymous):

no, i m BS(Hons) Physics student

OpenStudy (anonymous):

okay kool...